Trace this code

show the work

Public static int funcQ1(int m, int n) {

if (m < n)

return 0;

else return 1 + funQ1(m-n, n); }

• What is the value of funcQ1(6, 3), based on the code above?

• What is the value of funcQ1(9, 7), based on the code above?

• What is the value of funcQ1(-5, 1), based on the code above?
Answer #1

1)
funcQ1(6, 3)
= 1 + funcQ1(3, 3)
= 1 + 1 + funcQ1(0, 3)
= 1 + 1 + 0
= 2
Answer: 2

2)
funcQ1(9, 7)
= 1 + funcQ1(2, 7)
= 1 + 0
= 1
Answer: 1

1)
funcQ1(-5, 1)
= 0 (because -5 < 1)
Answer: 0